# Single attosecond XUV pulse source via light-wave controlled relativistic laser-plasma interaction: Thomson Back Scattering Scheme

### v1

@@ III Generation of Single RES by controlling driver waveform
Here we undertake a detailed study of the influence of light-wave control on RES formation by fine-tuning the electric field of the drive laser with the experimentally accessible parameters ($P,\Delta\phi$). Figure 2(a-b) summarizes the outcomes of 407 simulations in the ($P,\Delta\phi$)-space, where each run corresponds to one synthesized driver waveform spanning from single color $1\omega$ case through double color ($1\omega+2\omega$) to single color $2\omega$. The map on phase difference ranges from $-\pi$ to $\pi$, which is the periodicity of the total drive field.
@@
At a constant $P$, as the phase difference between two frequency components of the synthesized driving field varies from zero, the steepness of the waveform will change. While converting more energy to the second harmonic will modify the temporal width of the modulated part of the drive field once the phase difference is kept the same. The reason of obtaining high energy and dense electron beam at this parameters can be attributed to the shape of the laser waveform interacting with the target. Adding a particular portion of second harmonic to the fundamental beam modulates the waveform of the laser field, ensuring proper coupling and sufficient push force that leads to blow electrons out from the plasma foil under a specific condition.
@@
The role played by the longitudinal restoring force due to the space charge separation is crucial for dual-color drive beam particularly when the plasma electrons are expanded to the left hand side of the plasma ion slab, then being pushed to the forward direction, such that electrons within the left end of the bunch undergo larger amount of push force (electrostatic plus light pressure) and therefore can catch the slower electrons in the right side of the bunch. The electrostatic force accelerates them to relativistic velocity, the moment after which the RES is injected into the drive pulse field and experiences an unipolar field. The trapped bunch in the transmitted laser field is able to gain constant acceleration.
In order to understand the mechanism better, in Figure 2(g) and (h), we compare the temporal evolution of longitudinal and transverse components of velocity of the electrons involved in the interaction with single (black line) and two-color (magenta line) drive pulses from simulation time $10$ to $11.5T_{L}$, when the final stage of RES formation and acceleration occur. In case of $1\omega$ laser beam the electrons longitudinal velocity $\beta_{x}$ (black line in Figure 2(g)) reaches a peak in positive direction and then drops, every half cycle, since the laser push force is not sufficient for complete blowout. In Fig. 2(h) the transverse component of the electrons $\beta_{y}$ (magenta line) under the impact of two-color field peaks at $t=10.4T_{L}$, where the persistent laser magnetic field experienced by the electrons within the bunch exerts longitudinal force via $\beta_{y}\times B_{z}$. In case of two color drive pulse, both phases of expelling and accelerating of electrons take place within nearly half laser cycle, during which field does not change sign from $t=10.1T_{L}$ to $11.7T_{L}$. This field waveform with steeper slope not only avoids any additional electron oscillation, but also imposes an increasingly push force on bunched electrons, leading to synchronously reach relativistic velocity along positive x-axis (magenta line in Figure 2(g)). As the result of this larger longitudinal velocity of the RES and its lower energy spread, the temporal duration of the scattered source pulse is reduced by $3$ fold.
@@
In order to gain more insight on the evolution of RES properties we plot in Figure 2(i-j) electron bunch peak density and RES width for two-color Gaussian driver for the case of ($p=0.4,\Delta\phi=0$). The peak electron density can go upto $20n_{c}$ as seen in Figure 2(i). Due to the creation of highly dense RES, the intensity of the reflected source could be enhanced by more than one order of magnitude, as we have already presented in Figure 1(e). Figure 2(j) shows that the generated RES is sub-$10nm$ in thickness and implies that the bunch preserve its thickness during interaction with the source pulse. For the case of two-color beam, at $\sim 12.8T_{L}$ the intensity of the driving pulse drops which causes the electron bunch repulsion force to grow, therefore the RES density decreases to around $2n_{c}$. However, then by increasing the drive field value, the RES is again piled up.
@@ IV Control of attosecond pulse properties in single RES regime
Figure 3: The attosecond back-scattered pulse generated by in phase two-color drive beam with energy mixture ratio of $P=0.4$. The delay-amplitude maps for a $4$fs Gaussian source beam at a given amplitude $a_{0}$. (a) and (b) present scaling for Thomson back-scattered (TBS) pulse peak amplitude ratios $\eta=a_{0}^{TBS}/a_{0}$ and temporal compression $\xi=\tau/\tau_{{}_{TBS}}$ compared to the counter-propagating source pulse with amplitude $a_{0}$ and duration $\tau=4$fs. Temporal duration of each Thomson back-scattered field overlaid with its central frequency for the case of source pulses $a_{0}=0.2$ (c) and $a_{0}=2$ (b) are depicted. The TBS spectra consist of a peak (central) frequency with a FWHM bandwidth, indicated by a vertical bar. The bars are determined by fitting a Gaussian function on each spectrum. In each case of the map, the FWHM is calculated by fitting a Gaussian function on the squared field and spectrum.
@@
In the single RES regime achievable through light-wave control and investigated in the previous section, we notice that representative features that characterize the accelerated electron sheet (thickness, mean energy and peak electron density etc) evolves with time. This implies that two crucial aspects that are of utmost importance for isolated as pulse generation through TBS, are the instant at which the back-propagating ultrashort pulse starts interacting with the RES and the evolution of the RES within the interaction window. Thus, the delay between interaction of drive and source beams with the foil is of prime importance, and affects on the features of the scattered radiation. Another parameter that is of importance is the peak electric field of the incident source pulse represented by normalized amplitude $a_{0}$. Hence, we utilize this delay and $a_{0}$ as two control parameters for optimizing the properties of the generated attosecond pulse. As presented in Figure 2, a dual-color drive pulse with energy ratio parameter of $P=0.4$ and $\delta\phi=0$ is able to eject plasma electrons coherently to transmitted direction of the foil, producing a RES suitable for Thomson scattering of a counter-streaming source pulse. Without loss of generality, we investigate the as pulse characteristics generated from this RES, as a scenario representative of the single RES regime presented in Figure 2(a-b).
@@
Figure 3 shows the parametric scan of the features of reflected attosecond burst with respect to delay and amplitude of the Gaussian $\tau=4$fs source pulse. Here, the zero delay denotes the time, at which both the dive and source pulses arrive at the position of foil target at the same time ($t=6T_{L}$ in simulation time). In Figure 3(a) the ratio of peak amplitude of scattered AP to that of incident source $\eta$ is presented in colormap as a function of delay and $a_{0}$. Panel (b) in Figure 3 illustrates the achieved temporal compression value $\xi$ of the scattered pulse with respect to the $\tau=4$fs source pulse with variation in delay at different $a_{0}$. We observe that both $\eta$ and $\xi$ has strong dependence on delay and varies weakly with $a_{0}$. The highest compression ratio and field enhancement can be achieved at shorter delays for a constant $a_{0}$. Figure 3(a) shows that for $a_{0}<1$ (sub-relativistic intensities) $\eta$ shows a similar dependence on delay. As the source pulse strength is increased beyond the relativistic limit $a_{0}>1$ the enhancement factor $\eta$ reduces at the same delay, i.e for the same RES properties, indicating that the RES starts to be perturbed at such source pulse intensities. To get a deeper insight, we look into the AP properties as a function of delay at a non-relativistic ($a_{0}=0.2$) and relativistic ($a_{0}=2$) source pulse intensities.
Figure 4: Geometry of the dual-color Gaussian drive beam (represented by electric field $Ey$ proceeding to the right) and counter-propagating source beam (with electric field $E_{z}$ propagating to the left) just before the interaction at $1T_{L}$ in (a). Fields magnitude are normalized to $mc\omega_{L}/e$. The dynamics of the interaction is shown (b)-(d) where the single RES is accelerated in the forward direction and reflects a fraction of the source pulse marked by a rectangular box in (d). Panels (f)-(h) depict the on-axis spatial lineout of the drive pulse (green lines) overlaid with the electron density (gray lines) at $y=0$ averaged over $7$ adjacent grid cells. The distribution of the the single RES in transverse direction at $5T_{L}$ (black line) is demonstrated in (e) overlaid with the normalized transverse profile of the source pulse (light blue shade) and the initial transverse profile of the drive pulse on target in light green shade. Spatial, temporal and spectral features of the back-scattered source pulse reflected by the single RES generated by two-color Gaussian drive beam at simulation time $10T_{L}$ are depicted in panels (j)-(l), respectively. The temporal profile of the squared field is shown in (k) with $180$as duration at its FWHM, and the corresponding modulated spectrum in (l) without performing any frequency filtering. Maintaining all the interaction parameters, the radial size dependence of the generated single RES on the focal spot of the drive laser is plotted in panel (i). This analysis is performed at the moment $5T_{L}$ when the source pulse is supposed to interact with the RES. The green color coded dot represents the case in panel (e).
V Formation of single RES and as pulse generation with finite focal spot size
It is worth mentioning that the laser pulses employed experimentally access the proposed scheme are available in laser facilities such as ELI [50, 51, 3, 52] or elsewhere [1]. In addition, the effectiveness of replenishable sub-micrometer fluid targets in supporting relativistic laser-plasma interactions has been convincingly demonstrated in recent studies, particularly in the contexts of high harmonic generation [9] and laser-driven ion acceleration [53]. These targets offer a unique combination of ultra-thin geometry, high repetition rate compatibility, and surface smoothness, making them ideal for precision-controlled interactions with intense laser fields. Notably, chip-based microfluidic target platforms have achieved thicknesses below 50 nm [54], pushing the limits of target engineering for ultrafast science. When coupled with state-of-the-art, waveform-controlled laser systems, such ultra-thin fluid targets provide an ideal testbed for exploring light-wave-driven electron dynamics—a central theme of the present investigation. The ability to tailor the temporal and spatial structure of the laser field at sub-cycle resolution opens new avenues for coherent control of plasma surfaces, efficient RES formation, and attosecond-scale electron bunch generation.
In order to pinpoint not only the temporal and spectral characteristics of the scattered pulse but also its spatial features we have performed two-dimensional PIC code using fully relativistic electromagnetic PIC code, PICCANTE [55] with finite and realistic laser focal spot size. The simulation box which is from $-30\lambda_{L}$ to $30\lambda_{L}$ (in both $x$ and $y$ direction) consists of $60000\times 12000$ cells, in laser propagation direction and its normal, respectively. In each cell we have employed $60$ macro-electrons and $10$ macro-ions of fully ionized Carbon. The linearly polarized drive beam with electric field $E_{y}$ and central wavelength $\lambda_{L}=800$ propagates from left to right and impinges the pre-ionized $40$nm plasma foil located at $x=0$. Considering transverse Gaussian profile of the intensity $\propto\exp(-4\ln{2}y^{2}/w_{0}^{2})$ ), the beam spot size at target position is $w_{0}=18$µm (in FWHM of intensity), as well as a Gaussian temporal profile of duration $10$fs at FWHM of intensity envelope. This multi-cycle drive beam with normalized field strength of $a_{0}^{d}=20$ is temporally and spatially combined with its second harmonic with a ratio of $P=0.4$ which resembles the optimal drive waveform obtained from 2D plane wave simulations presented previously.
@@ V Formation of single RES and as pulse generation with finite focal spot size
Figure 4(a) depicts the configuration of the interaction where the intense dual-color drive beam reaches the foil position at $1T_{L}$ simulation time. The initial electron density of the plasma slab is set to be $32n_{c}$, which is slightly greater than of that in $2$D plane wave simulations ($21n_{c}$). This is due to the fact that the relativistic ponderomotive force associated with the focused drive pulse introduces a transverse scattering push on the plasma electrons and expels them to the region, at which light intensity is lower. Therefore, the effective electron density of the plasma, and consequently the electrostatic field at the interaction surface will be reduced. This effect is thus compensated by increasing the initial electron density of the target to maintain the proper interaction conditions for the generation of single RES. The focused drive beam expels the plasma electrons and accelerates them in phase along its propagation direction. After a delay of $3T_{L}$, since when the leading edge of the drive has reached the foil front position, a low intensity counter-propagating source pulse with cross-polarized electric field $E_{z}$ interacts with the driven RES, which travels in positive $x$ direction. The tightly focused Gaussian source beam has $4$fs temporal duration and $2.9\lambda_{L}$ transverse dimension in FWHM of its envelope on target at $x=0$. At this instant of interaction, the plasma slab which is subject to intense electromagnetic field of the drive beam turns relativistically transparent, leading to the transmission of drive pulse. The temporal evolution of the single RES generation process is shown in Figure 4(b)-(d), highlighting the reflection of a portion of the source pulse by the RES. The transverse extent of the generated single RES is illustrated in Figure. 4(d) in black line. Figure. 4(d) shows that the source pulse is covered by the size of the RES formed. The lineout at $y=0$ (averaged over $0.2\lambda_{L}$) of the drive beam field, which is plotted in Figure 4(f) to (h) in green curve, substantiate the fact that analogous to previous results, plasma electrons are ejected forward by the same cycle of the drive pulse. Afterward, the accelerated RES and drive pulse propagate together in forward direction. Since, the drive pulse is loosely focused, the co-moving RES undergoes negligible amount of ponderomotive scattering force, maintaining its structure for a longer time.
@@
As a consequence of interaction of this RES and counter-propagating source beam a single sub fs radiation, consisting of high frequencies is generated and propagates in forward direction with polarization similar to that of source beam. The scattered portion of the source pulse, which advances to the positive $x$-axis is illustrated in Figure 4(j-l). The spatial dimension of this attosecond burst in Figure 4(j) also corroborates the generation of a single RES during interaction of a dual-color Gaussian drive pulse. Panel (k) plots the envelope of the squared field with temporal duration of $180$as, which is relatively consistent previous results. The structured spectrum corresponding to lineout of $E_{z}$ (averaged over $0.1\lambda_{L}$) is shown in Figure 4(l), comprising a central peak at $10\omega_{L}$ and a low energy portion extending up to $20\omega_{L}$.

### v2

@@ III Generation of Single RES by controlling driver waveform
Here we undertake a detailed study of the influence of light-wave control on RES formation by fine-tuning the electric field of the drive laser with the experimentally accessible parameters ($P,\Delta\phi$). Figure 2(a-b) summarizes the outcomes of 407 simulations in the ($P,\Delta\phi$)-space, where each run corresponds to one synthesized driver waveform spanning from single color $1\omega$ case through double color ($1\omega+2\omega$) to single color $2\omega$. The map on phase difference ranges from $\pi$ to $\pi$, which is the periodicity of the total drive field.
@@
At a constant $P$, as the phase difference between two frequency components of the synthesized driving field varies from zero, the steepness of the waveform will change. While converting more energy to the second harmonic will modify the temporal width of the modulated part of the drive field once the phase difference is kept the same. Thus, the reason of obtaining high energy and dense electron beam at this parameters can be attributed to the shape of the laser waveform interacting with the target. Adding a particular portion of second harmonic to the fundamental beam modulates the waveform of the laser field, ensuring proper coupling and sufficient push force that leads to blow electrons out from the plasma foil under a specific condition.
@@
The role played by the longitudinal restoring force due to the space charge separation is crucial for dual-color drive beam particularly when the plasma electrons are expanded to the left hand side of the plasma ion slab, then being pushed to the forward direction, such that electrons within the left end of the bunch undergo larger amount of push force (electrostatic plus light pressure) and therefore can catch the slower electrons in the right side of the bunch. The electrostatic force accelerates them to relativistic velocity, the moment after which the RES is injected into the drive pulse field and experiences an unipolar field. The trapped bunch in the transmitted laser field is able to gain energy until it dephases [51, 52].
To understand the mechanism better, in Figures 2(g) and 2(h), we compare the temporal evolution of longitudinal and transverse components of velocity of the electrons involved in the interaction with single (black line) and two-color (magenta line) drive pulses from simulation time $10$ to $11.5T_{L}$, when the final stage of RES formation and acceleration occur. In case of $1\omega$ laser beam the electrons longitudinal velocity $\beta_{x}$ (black line in Figure 2(g)) reaches a peak in positive direction and then drops, every half cycle, since the laser push force is not sufficient for complete blowout. In Figure 2(h) the transverse component $\beta_{y}$ (magenta line), of the electron velocity, under the mpact of two-color field peaks at $t=10.4T_{L}$, where the persistent laser magnetic field experienced by the electrons within the bunch exerts a longitudinal force via $\sim\beta_{y}B_{z}\hat{x}$. In case of two color drive pulse, both phases of expelling and accelerating of electrons take place within nearly half laser cycle, during which field does not change sign from $t=10.1T_{L}$ to $11.7T_{L}$. This field waveform with steeper slope not only avoids any additional electron oscillation, but also imposes an increasingly push force on bunched electrons, leading to synchronously reach relativistic velocity along positive x-axis (magenta line in Figure 2(g)). As the result of this larger longitudinal velocity of the RES and its lower energy spread, the temporal duration of the scattered source pulse is reduced by $3$ fold.
Figure 3: The attosecond back-scattered pulse generated by in phase two-color drive beam with energy mixture ratio of $P=0.4$. The delay-amplitude maps for a 44fs Gaussian source beam at a given amplitude $a_{0}$. (a) and (b) present scaling for Thomson back-scattered (TBS) pulse peak amplitude ratios $\eta=a_{0}^{TBS}/a_{0}$ and temporal compression $\xi=\tau/\tau_{{}_{TBS}}$ compared to the counter-propagating source pulse with amplitude $a_{0}$ and duration $\tau=4$fs. Temporal duration of each Thomson back-scattered field overlaid with its central frequency for the case of source pulses $a_{0}=0.2$ (c) and $a_{0}=2$ (b) are depicted. The TBS spectra consist of a peak (central) frequency with a FWHM bandwidth, indicated by a vertical bar. In each case, the FWHM is calculated by fitting a Gaussian function on the squared field and spectrum.
@@
In order to gain more insights into the evolution of RES properties we plot in Figure 2(i-j) electron bunch peak density and RES width for two-color Gaussian driver for the case of ($P=0.4,\Delta\phi=0$). The peak electron density can go upto $20n_{c}$ as seen in Figure 2(i). Due to the creation of highly dense RES, the intensity of the reflected source could be enhanced by more than one order of magnitude, as we have already presented in Figure 1(e). Figure 2(j) shows that the generated RES is sub-$10$ nm in thickness and implies that the bunch preserve its thickness during interaction with the source pulse. For the case of two-color beam, at $\sim 12.8T_{L}$ the intensity of the driving pulse drops which causes the electron bunch repulsion force to grow, therefore the RES density decreases to around $2n_{c}$. However, then by increasing the drive field value, the RES is again piled up.
@@ IV Control of attosecond pulse properties in single RES regime
In the single RES regime achievable through light-wave control and investigated in the previous section, we notice that representative features that characterize the accelerated electron sheet (thickness, mean energy and peak electron density etc) evolve with time. This implies that two crucial aspects that are of utmost importance for isolated as pulse generation through TBS, are the instant at which the back-propagating ultrashort pulse starts interacting with the RES and the evolution of the RES within the interaction window. Thus, the delay between interaction of drive and source beams with the foil is of prime importance, and affects on the features of the scattered radiation. Another parameter that is of importance is the peak electric field of the incident source pulse represented by normalized amplitude $a_{0}$. Hence, we utilize this delay and $a_{0}$ as two control parameters for optimizing the properties of the generated attosecond pulse. As presented in Figure 2, a dual-color drive pulse with energy ratio parameter of $P=0.4$ and $\Delta\phi=0$ is able to eject plasma electrons coherently to transmitted direction of the foil, producing a RES suitable for Thomson scattering of a counter-streaming source pulse. Without loss of generality, we investigate the AP characteristics generated from this RES, as a scenario representative of the single RES regime presented in Figure 2(a-b).
The accelerated RES is irradiated by a counter propagating source pulse, having a crossed linear polarization along z direction $(E_{z})$. The backscattered field is an isolated attosecond pulse in of 110asas duration as shown in Figure  1(e). For representing the AP properties we have defined two parameters. The generated as pulse field enhancement is captured through peak amplitude ratios $\eta=a_{0}^{TBS}/a_{0}$, where $a_{0}$ is the normalized field amplitude of the incident source pulse and $a_{0}^{TBS}$ is the same for the as pulse. We also define temporal compression $\xi=\tau/\tau_{{}_{TBS}}$, where $\tau=4$fs is the duration of the Gaussian source pulse and $\tau_{{}_{TBS}}$ is the pulse duration of the generated as pulses.
Figure 3 shows the parametric scan of the features of reflected attosecond burst with respect to delay and amplitude of the Gaussian $\tau=4$fs source pulse. Here, the zero delay denotes the time, at which both the drive and source pulses arrive at the position of foil target at the same time ($t=6T_{L}$ in simulation time). In Figure 3(a) the ratio of peak amplitude of scattered AP to that of incident source $\eta$ is presented in colormap as a function of delay and $a_{0}$. Panel (b) in Figure 3 illustrates the achieved temporal compression value $\xi$ of the scattered pulse with respect to the $\tau=4$fs source pulse with variation in delay at different $a_{0}$. We observe that both $\eta$ and $\xi$ has strong dependence on delay and varies weakly with $a_{0}$. The highest compression ratio and field enhancement can be achieved at shorter delays for a constant $a_{0}$. Figure 3(a) shows that for $a_{0}<1$ (sub-relativistic intensities) $\eta$ shows a similar dependence on delay. As the source pulse strength is increased beyond the relativistic limit $a_{0}>1$ the enhancement factor $\eta$ reduces at the same delay, i.e for the same RES properties, indicating that the RES starts to be perturbed at such source pulse intensities. To get a deeper insight, we look into the AP properties as a function of delay at a non-relativistic ($a_{0}=0.2$) and relativistic ($a_{0}=2$) source pulse intensities.
Figure 4: Geometry of the dual-color Gaussian drive beam (represented by electric field EyEy proceeding to the right) and counter-propagating source beam (with electric field $E_{z}$ propagating to the left) just before the interaction at $1T_{L}$ in (a). Electric field magnitudes are normalized to $mc\omega_{L}/e$. The dynamics of the interaction is shown (b)-(d) where the single RES is accelerated in the forward direction and reflects a fraction of the source pulse marked by a rectangular box in (d). Panels (f)-(h) depict the on-axis spatial lineout of the drive pulse (green lines) overlaid with the electron density (gray lines) at $y=0$ averaged over 77 adjacent grid cells. The distribution of the the single RES in transverse direction at $5T_{L}$ (black line) is demonstrated in (e) overlaid with the normalized transverse profile of the source pulse (light blue shade) and the initial transverse profile of the drive pulse on target in light green shade. Spatial, temporal and spectral features of the back-scattered source pulse reflected by the single RES generated by two-color Gaussian drive beam at simulation time $10T_{L}$ are depicted in panels (j)-(l), respectively. The temporal profile of the squared field is shown in (k) with 180180as duration at its FWHM, and the corresponding modulated spectrum in (l) without performing any frequency filtering. Maintaining all the interaction parameters, the radial size dependence of the generated single RES on the focal spot of the drive laser is plotted in panel (i). This analysis is performed at the moment $5T_{L}$ when the source pulse is supposed to interact with the RES. The green color coded dot represents the case in panel (e).
V Formation of single RES and AP generation with finite focal spot size
@@
Figure 4(a) depicts the configuration of the interaction where the intense dual-color drive beam reaches the foil position at $1T_{L}$ simulation time. The initial electron density of the plasma slab is set to be $32n_{c}$, which is slightly greater than of that in 22D plane wave simulations ($21n_{c}$). This is due to the fact that the relativistic ponderomotive force associated with the focused drive pulse introduces a transverse scattering push on the plasma electrons and expels them to the region, at which light intensity is lower. Therefore, the effective electron density of the plasma, and consequently the electrostatic field at the interaction surface will be reduced. This effect is thus compensated by increasing the initial electron density of the target to maintain the proper interaction conditions for the generation of single RES. The focused drive beam expels the plasma electrons and accelerates them in phase along its propagation direction. After a delay of $3T_{L}$, since when the leading edge of the drive has reached the foil front position, a low intensity counter-propagating source pulse with cross-polarized electric field $E_{z}$ interacts with the driven RES, which travels in positive $x$ direction. The tightly focused Gaussian source beam has 44fs temporal duration and $2.9\lambda_{L}$ transverse dimension in FWHM of its envelope on target at $x=0$. At this instant of interaction, the plasma slab which is subject to intense electromagnetic field of the drive beam turns relativistically transparent, leading to the transmission of drive pulse. The temporal evolution of the single RES generation process is shown in Figure 4(b)-(d), highlighting the reflection of a portion of the source pulse by the RES. The transverse extent of the generated single RES is illustrated in Figure. 4(d) in black line. Figure. 4(d) shows that the source pulse is covered by the size of the RES formed. The lineout at $y=0$ (averaged over $0.2\lambda_{L}$) of the drive beam field $E_{y}$, which is plotted in Figure 4(f) to 4(h) in green curve, substantiate the fact that analogous to previous results, plasma electrons are ejected forward by the same cycle of the drive pulse. Afterward, the accelerated RES and drive pulse propagate together in forward direction. Since, the drive pulse is loosely focused, the co-moving RES undergoes negligible amount of ponderomotive scattering force, maintaining its structure for a longer time.
@@
As a consequence of interaction of this RES and counter-propagating source beam a single sub fs radiation, consisting of high frequencies is generated and propagates in forward direction with polarization similar to that of source beam. The scattered portion of the source pulse, which advances to the positive $x$-axis is illustrated in Figure 4(j-l). The spatial dimension of this attosecond burst in Figure 4(j) also corroborates the generation of a single RES during interaction of a dual-color Gaussian drive pulse. Panel 4(k) plots the envelope of the squared field with temporal duration of 180180as, which is qualitatively consistent with the previous results. The structured spectrum corresponding to lineout of $E_{z}$ (averaged over $0.1\lambda_{L}$) is shown in Figure 4(l), comprising a central peak at $10\omega_{L}$ and a low energy portion extending up to $20\omega_{L}$.